The equation 0.01x + 4= 14 has solution?


1
Expert's answer
2021-08-31T04:49:03-0400

"0.01^x+4=14;"

"0.01^x=14-4;"

"0.01^x=10"

Taking logarithm of both sides, we get:

"log(0.01^x)=log(10);"

"x\times{log(0.01)}=log(10);"

"x=\frac{log(10)}{log(0.01)}=\frac{1}{-2}=-0.5"


Answer: yes, there is a solution, and x = -0.5